What is the sum of the first 45 terms of the following sequence?
20, 18, 16, . . .

Answer by stanbon(75887)

What is the sum of the first 45 terms of the following sequence?
20, 18, 16, . . .
----
a = 20
d = -2
---
S(n) = (n/2)(a+a(n))
----
a(n) = a + (n-1)d
a(45) = 20 + 44*-2 = 20-88 = -68
-------------------------------------
S(45) = (45/2)(20 - 68)
---
= 22.5(-48)
---
= -1080
===============
Cheers,
Stan H.
===============